### v2

# Friendship-paradox paradox: Do most people’s friends really have more friends than they do?

Abstract
The classical friendship paradox asserts that, on average, an individual’s neighbors have a higher degree than the individual. This statement concerns network-level means and does not describe how often a typical node is locally dominated by its neighbors. Motivated by this distinction, we develop a framework that separates mean-based friendship paradox inequalities from two majority-type quantities: a global fraction measuring how many nodes have a degree smaller than the mean degree of their neighbors, and a local fraction based on hub centrality that measures how many nodes are dominated in a median-based sense. We show that neither fraction is constrained by the classical friendship paradox and that they can behave independently of each other. A simple example and two empirical networks illustrate how quadrant patterns in the joint distribution of a node’s degree and its neighbors’ degree determine the signs and magnitudes of the two fractions, and how left- or right-skewed degree distributions of neighboring nodes can yield opposite conclusions for mean-based and median-based comparisons. The resulting framework offers a clearer distinction between population averages and local majority relations and provides a foundation for future analyses of local advantage, disadvantage, and perception asymmetry in complex networks.
I Introduction
The friendship paradox (FP), first articulated in Ref. [1], is usually summarized as the claim that “your friends have more friends than you do.” This phrase has become a standard way to describe the bias that arises when individuals compare their own connectivity with that of their neighbors in a network. Behind this simple statement, however, lie several different ways to formalize how one samples nodes and how one performs the comparison. At the level of network-wide averages, the classical FP admits two natural formulations. The first, sometimes called the alter-based version, focuses on nodes reached by following randomly chosen edges and compares their average degree with the population average. The second, the ego-based version, averages for each node the degrees of its neighbors and then compares that average with the mean degree over nodes. These two formulations coincide in degree-uncorrelated networks but can differ when degree-degree correlations are present. Recent work [2, 3] has clarified how their difference is controlled by correlations between a node’s degree and the average degree of its neighbors.
A separate but closely related family of questions asks not about averages over the whole network, but about the fraction of nodes that are locally disadvantaged relative to their neighbors. One global majority-type criterion asks whether more than half of the nodes have a degree smaller than the mean degree of their neighbors. A different, more local majority criterion uses the hub centrality [4], or the fraction-based peer pressure [5], which measures for each node the fraction of neighbors whose degree is smaller than its own. In that framework, one can ask how often a node is outnumbered by higher-degree neighbors in a median sense. These majority-type inequalities are conceptually different from the original FP because they concern medians or fractions of nodes rather than network-level means.
In practice, the colloquial phrasing of the friendship paradox is often taken to express a majority claim such as “for most people, their friends have more friends than they do.” This interpretation presumes that a typical individual is locally less connected than the people around them. The classical paradox, however, makes no such statement. It concerns only comparisons of network-wide averages and offers no guarantee about what happens for most nodes, or even for any specific node. As the examples in this paper illustrate, it is entirely possible for the paradox to hold in its traditional, mean-based form while a majority of nodes are in fact more connected than their neighbors on average. Conversely, one may also find networks where many nodes appear locally disadvantaged despite the population averages suggesting otherwise. These discrepancies arise because local comparisons depend sensitively on the distribution of neighbor degrees around each node, not merely on global averages. This observation motivates the need to distinguish clearly between population-level statements and majority-type, node-level assessments, and it sets the stage for the systematic framework developed in the remainder of the paper.
The purpose of this work is to separate these notions clearly and to provide a minimal mathematical language for doing so. After introducing the basic notation in the next section, we review the two classical FP inequalities and then define two majority-type quantities: a global mean-based fraction that encodes the question “is it likely that your friends have more friends than you?” and a local median-based fraction, based on hub centrality, that encodes the question “is it likely that most of your neighbors have more friends than you?” We then analyze how these quantities behave in a toy example and in two empirical networks, highlighting the structural reasons why they can be different. The main goal is to offer a conceptually transparent framework that can support future quantitative work on local advantage, disadvantage, and perception asymmetry in complex networks.
II Mathematical framework
II.1 Network notation
Consider a simple undirected network with the adjacency matrix $A=\{a_{ij}\}$, where $a_{ij}=1$ if nodes $i$ and $j$ are connected and $a_{ij}=0$ otherwise. The degree of node $i$ is
$k_{i}=\sum_{j}a_{ij}\,$, (1)
and the neighbor set and neighbor-degree multiset are
$\mathcal{N}(i)=\{j:a_{ij}=1\}\,,\qquad\mathcal{K}_{i}=\{k_{j}:j\in\mathcal{N}(i)\}\,.$ (2)
The mean neighbor degree of node $i$ is
$k_{nn}(i)=\frac{1}{k_{i}}\sum_{j\in\mathcal{N}(i)}k_{j}\,$, (3)
defined whenever $k_{i}>0$. Averages over nodes with respect to the empirical degree distribution $p(k)$ are written as
$\langle f(k)\rangle_{\mathrm{n}}=\sum_{k}f(k)\,p(k)\,$, (4)
where the notational convention $\langle\cdots\rangle_{\mathrm{n}}$ follows Ref. [3].
II.2 Classical alter-based and ego-based FP inequalities
The alter-based FP considers the degree of a node reached by following a uniformly random edge. The corresponding mean degree,
$\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}=\frac{\langle k^{2}\rangle_{\mathrm{n}}}{\langle k\rangle_{\mathrm{n}}}\geq\langle k\rangle_{\mathrm{n}}\,$, (5)
The ego-based FP instead averages the per-node mean neighbor degree:
$\langle k_{nn}\rangle_{\mathrm{n}}=\frac{1}{N}\sum_{i}k_{nn}(i)\geq\langle k\rangle_{\mathrm{n}}\,.$ (6)
Those are the original formulations presented in Ref. [1]. Recent works [2, 3] have shown that the difference between the left-hand sides of Eqs. (5) and (6) is determined by the covariance between $k_{i}$ and $k_{nn}(i)$ over nodes [3], i.e.,
$\langle k_{\text{friend}}\rangle_{\mathrm{n}}-\langle k_{nn}\rangle_{\mathrm{n}}=\frac{1}{\langle k\rangle_{\mathrm{n}}}\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})\,$, (7)
(verified in Table 1 for the example networks analyzed in this work) or equivalently (the equivalence is proved in Ref. [3]) as presented in Ref. [2],
$\langle k_{nn}\rangle_{\mathrm{n}}-\langle k_{\text{friend}}\rangle_{\mathrm{n}}=\rho\sqrt{\left(\frac{\kappa_{1}\kappa_{3}-\kappa_{2}^{2}}{\kappa_{1}}\right)\left(\kappa_{-1}-\kappa_{1}^{-1}\right)}\,$, (8)
where $\rho=\mathrm{Corr_{\mathrm{e}}}\left(D_{D},D_{O}^{-1}\right)$, which is the Pearson correlation coefficient between the degree $D_{D}$ of the origin node (“the origin degree” in Ref. [2]) and the reciprocal of the degree $D_{O}$ of the destination node (“the inverse destination degree” in Ref. [2]) and $\kappa_{m}=\left(\sum_{i=1}^{N}k_{i}^{m}\right)/N$ is the $m$th moment of degrees.
Note that both inequalities in Eqs. (5) and (6) are statements about averages across the network and do not refer to the fraction of nodes that are individually dominated by their neighbors.
II.3 Global mean-based majority fraction
To formalize a global majority-type interpretation, define
$\phi_{\mathrm{global}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{k_{i}<k_{nn}(i)\}}\,$, (9)
where $\mathbf{1}_{\{\cdot\}}$ denotes the indicator function, equal to $1$ if the condition inside the braces is true and 0 otherwise. The quantity $\phi_{\mathrm{global}}$ represents the fraction of nodes whose degree is smaller than the mean degree of their neighbors. The colloquial claim “it is likely that your friends have more friends than you” corresponds to the majority condition
$\phi_{\mathrm{global}}>\frac{1}{2}\,.$ (10)
Crucially, the inequalities in Eqs. (5) and (6) do not imply any constraint on $\phi_{\mathrm{global}}$. Both $\phi_{\mathrm{global}}>1/2$ and $\phi_{\mathrm{global}}<1/2$ are possible, even in networks where the classical FP holds in both alter-based and ego-based forms. The AFB network example discussed later provides an explicit case where $\phi_{\mathrm{global}}<1/2$ despite the FP being satisfied.
II.4 Local median-based majority fraction
The median-based perspective introduced in the context of local hubs [4] is encoded by the hub centrality
$h_{i}=\frac{1}{k_{i}}\sum_{j\in\mathcal{N}(i)}\mathbf{1}_{\{k_{j}<k_{i}\}}\,$, (11)
which measures the fraction of neighbors whose degree is smaller than that of node $i$. This quantity is closely related to the neighborhood median, but the relation requires some care. When neighbor degrees are treated as a multiset $\mathcal{K}_{i}$, the condition
$h_{i}<\frac{1}{2}\,$, (12)
implies that fewer than half of the neighbors have degree strictly smaller than $k_{i}$. This ensures that at least half of the neighbors have degree $\geq k_{i}$, which is consistent with the inequality
$k_{i}\leq\mathcal{K}_{i}^{(1/2)}\,$, (13)
where $\mathcal{K}_{i}^{(1/2)}$ denotes the median of the multiset $\mathcal{K}_{i}$. In the absence of ties and for sufficiently large neighborhoods, this condition effectively coincides with the strict comparison
$k_{i}<\mathcal{K}_{i}^{(1/2)}\,$, (14)
and the two formulations become asymptotically equivalent in the thermodynamic limit. However, for finite graphs with ties or very small neighborhoods, the equivalence does not in general hold as a biconditional. These boundary effects do not arise from the median formalism itself but from the strict inequality built into the definition of hub centrality, which treats equal-degree neighbors as neither “higher” nor “lower.” With these qualifications, the interpretation remains clear: a node with $h_{i}<1/2$ is locally dominated by higher-degree neighbors in a majority (median-based) sense, although the precise relationship with the median requires attention to ties and small-neighborhood effects.
The corresponding population fraction is
$\phi_{\mathrm{local}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{h_{i}<1/2\}}\,$, (15)
which answers the question “is it likely that most of your neighbors have more friends than you?”, i.e., whether
$\phi_{\mathrm{local}}>\frac{1}{2}\,$, (16)
holds or not. Neither $\phi_{\mathrm{local}}$ nor its comparison with one-half is constrained by the classical FP. As the toy example and the two empirical networks in later sections will show, $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ can take values on opposite sides of one-half and vary almost independently, even though both are derived from comparisons between a node and its neighbors.
The local median-based formulation in Eqs. (11) and (12) is mathematically equivalent to the “fraction-based peer pressure” measure introduced in Ref. [5], apart from the treatment of ties when $k_{i}=k_{j}$ (ties are excluded both in Eq. (11) and in Ref. [5]). A conceptual difference arises in how the node-level quantities are aggregated. Ref. [5] defines a network-level peer pressure by averaging the local pressures over all nodes, whereas Eq. (15) in the present work takes the fraction of nodes satisfying the median-based condition $h_{i}<1/2$. This contrast leads to an interesting duality: Eq. (9) represents a network-level fraction derived from local comparisons of mean values, while Ref. [5] constructs a network-level average derived from local comparisons of median values. Finally, Eq. (15) is a combination of both: it represents a network-level fraction derived from local comparisons of median values.
In the sections that follow, these quantities are evaluated in a simple toy network, in the ZKC network, and in the AFB network, and their joint behavior is interpreted in terms of their local organization of joint distributions of $h_{i}$ and $k_{nn}$.
III Systematic Difference Between the Criteria
The four quantities, $\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}$, $\langle k_{nn}\rangle_{\mathrm{n}}$, $\phi_{\mathrm{global}}$, and $\phi_{\mathrm{local}}$, encode two fundamentally different modes of comparison: mean-based relations at the level of population averages, and median-based relations expressed as majority conditions at the level of individual nodes. Although the classical FP guarantees the two mean-based inequalities, no analogous constraint binds the majority-type quantities. The mean and the median of the same neighbor-degree multiset $\mathcal{K}_{i}$ can respond quite differently to local heterogeneity. A single high-degree neighbor can substantially raise $k_{nn}(i)$ while leaving $\mathcal{K}_{i}^{(1/2)}$ almost unchanged, and degree-degree correlations can further amplify such discrepancies across the network.
Consequently, the two majority-type inequalities $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ need not align with each other, nor with the classical FP. Their values may lie above or below one-half depending on how degree values and neighborhood structures are distributed. This naturally motivates the question of how small or simple a network can be while already displaying such divergence. The following toy example provides a simple construction in which $\phi_{\mathrm{global}}<1/2$ and $\phi_{\mathrm{local}}<1/2$ hold simultaneously, even though the classical FP inequalities remain satisfied.
III.1 A toy example with $\phi_{\mathrm{global}}<1/2$ and $\phi_{\mathrm{local}}<1/2$
12345
Figure 1: A five-node toy network with $\phi_{\mathrm{global}}=\phi_{\mathrm{local}}=2/5<1/2$. Nodes $1$, 22, and 33 form a triangle, and nodes $4$ and $5$ are connected to all of 11, 22, and $3$. Degrees are $k_{1}=k_{2}=k_{3}=4$ and $k_{4}=k_{5}=3$. The high-degree nodes 1133 are not locally dominated by their neighbors, whereas the low-degree nodes $4$ and 55 are.
Figure 1 presents a five-node toy network that explicitly realizes the case where both $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ become less than $1/2$. In this network, the degrees are
$k_{1}=k_{2}=k_{3}=4\,,\qquad k_{4}=k_{5}=3\,.$ (17)
For nodes $1$, 22, and $3$, the neighbor degrees are $\{4,4,3,3\}$, so
$k_{nn}(1)=k_{nn}(2)=k_{nn}(3)=\frac{4+4+3+3}{4}=3.5<4\,$, (18)
and the mean-based inequality $k_{i}<k_{nn}(i)$ fails at these nodes. For nodes $4$ and $5$, the neighbor degrees are $\{4,4,4\}$, so
$k_{nn}(4)=k_{nn}(5)=\frac{4+4+4}{3}=4>3\,$, (19)
and they satisfy $k_{i}<k_{nn}(i)$. Hence
$\phi_{\mathrm{global}}=\frac{1}{5}\bigl(0+0+0+1+1\bigr)=\frac{2}{5}<\frac{1}{2}\,.$ (20)
For the median-based quantity, the hub centralities are
$h_{1}=h_{2}=h_{3}=\frac{2}{4}=\frac{1}{2}\,,\qquad h_{4}=h_{5}=\frac{0}{3}=0\,.$ (21)
Nodes $1$, 22, and $3$ therefore have exactly half of their neighbors with smaller degree and do not contribute to $\phi_{\mathrm{local}}$, whereas nodes $4$ and $5$ satisfy $h_{i}<1/2$ and are counted as being locally dominated in the median sense. Although the cases $h_{i}=1/2$ lie precisely at the decision boundary, $k_{i}=4$ is definitely a dominant one compared to the only other degree value $k_{i}=3$ in this example, so we retain them in this form for illustrative clarity. Consequently,
$\phi_{\mathrm{local}}=\frac{1}{5}\bigl(0+0+0+1+1\bigr)=\frac{2}{5}<\frac{1}{2}\,.$ (22)
This example shows that it is easy to construct situations where a minority of nodes are locally dominated (in either the mean- or median-based sense). The toy network highlights that $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ can take values below one-half, despite the classical FP inequalities.
IV Empirical Examples: Karate Club and American Football Networks
Table 1: Numerical values of the four quantities $\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}$, $\langle k_{nn}\rangle_{\mathrm{n}}$, $\phi_{\mathrm{global}}$, and $\phi_{\mathrm{local}}$ for the toy network in Fig. 1, the ZKC network [6], and the AFB network [7, 8]. The classical FP quantities are always larger than $\langle k\rangle_{\mathrm{n}}$, while the majority-type quantities $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ vary widely across networks, demonstrating their independence from mean-based FP conditions. The relation between the two classical FPs [3] in Eq. (7) is also confirmed for all of the cases.
Fig. 1 ZKC AFB
NN 55 $34$ $115$
$\langle k\rangle_{\mathrm{n}}$ $3.6$ $4.588235$ $10.660870$
$\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}$ $11/3\approx 3.67$ $7.769231$ $10.734095$
$\langle k_{nn}\rangle_{\mathrm{n}}$ $37/10=3.7$ $9.610210$ $10.721979$
$\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})/\langle k\rangle_{\mathrm{n}}$ $1/30\approx-0.03$ 1.840979-1.840979 $0.012116$
$\phi_{\mathrm{global}}$ $0.4$ $0.852941$ $0.434783$
$\phi_{\mathrm{local}}$ $0.4$ $0.705882$ $0.843478$
Figure 2: Density plots for the ZKC network [6], illustrating node-level relationships among degree $k_{i}$, mean neighbor degree $k_{nn}(i)$, hub centrality $h_{i}$, and the difference $k_{i}-k_{nn}(i)$ between a node’s degree and the average value of its neighbors’ degrees. The thresholds for the mean-based or median-based comparison, $k_{nn}(i)=k_{i}$ and $h_{i}=1/2$, are plotted as the white dashed lines. Panel (a) shows $k_{i}$ versus $k_{nn}(i)$, providing a node-level view of the ego-based FP; most points lie above the diagonal $k_{nn}(i)=k_{i}$, consistent with the large value $\phi_{\mathrm{global}}\approx 0.85$. Panel (b) shows $k_{i}$ versus $h_{i}$, emphasizing the median-based notion of local dominance. Panel (c) plots $h_{i}$ against $k_{i}-k_{nn}(i)$, revealing that nodes with negative mean-based contrast tend to have low hub centrality, indicating simultaneous mean-based and median-based disadvantage. Together, the panels illustrate why both $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ are high in this network.
Figure 3: Density plots for the AFB network [7, 8], illustrating the same three relationships shown in Fig. 2. Panel (a) displays $k_{i}$ versus $k_{nn}(i)$ and shows that a majority of points lie below the diagonal $k_{nn}(i)=k_{i}$, consistent with $\phi_{\mathrm{global}}\approx 0.43<1/2$, i.e., a majority of teams have degree larger than the mean degree of their neighbors, even though the classical FP in Eqs. (5) and (6) holds. Panel (b) shows $k_{i}$ versus $h_{i}$, and panel (c) shows $h_{i}$ versus $k_{i}-k_{nn}(i)$; both plots reveal that a large fraction of nodes satisfy $h_{i}<1/2$, yielding $\phi_{\mathrm{local}}\approx 0.84$. These contrasting patterns demonstrate that mean-based and median-based majority-type inequalities in Eqs. (10) and (16) can also be quite different, in contrast to the toy network in Fig. 1 with $\phi_{\mathrm{global}}=\phi_{\mathrm{local}}<1/2$.
To illustrate the four measures discussed in this paper, consider two simple (and perhaps overused, but not in this context) empirical data: Zachary’s karate club (ZKC) network [6] and the National Collegiate Athletic Association Division I American football (AFB) network compiled and published in Refs. [7, 8]. For each network, the two classical FP quantities $\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}$ and $\langle k_{nn}\rangle_{\mathrm{n}}$ are computed, along with the two median-based fractions defined earlier:
$\phi_{\mathrm{global}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{k_{i}<k_{nn}(i)\}}\,,\qquad\phi_{\mathrm{local}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{h_{i}<1/2\}}\,.$ (23)
Table 1 summarizes the numerical values of the four quantities for all three networks.
The density plots in Figs. 2 and 3 provide a finer-grained view of how mean-based and median-based comparisons diverge at the node level. In the ZKC network (Fig. 2), panel (a) shows that a large concentration of nodes lies above the diagonal $k_{nn}(i)=k_{i}$, consistent with $\phi_{\mathrm{global}}\approx 0.85$: most individuals have a degree smaller than the mean degree of their neighbors. Panel (b) demonstrates that many of these same nodes have low hub centrality $h_{i}$, and panel (c) confirms that negative values of $k_{i}-k_{nn}(i)$ tend to co-occur with low $h_{i}$. The three panels together indicate that, for this network, mean-based and median-based notions of being “less connected than one’s neighbors” align for most nodes, yielding simultaneously large values of $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$.
In the AFB network (Fig. 3), however, panel (a) shows a substantial density of nodes below the diagonal, consistent with $\phi_{\mathrm{global}}\approx 0.43<1/2$: the majority of teams have degree larger than the mean degree of their neighbors. Panels (b) and (c), however, reveal a contrasting pattern: most nodes satisfy $h_{i}<1/2$, producing $\phi_{\mathrm{local}}\approx 0.84$. Thus, even though a majority of teams exceed the mean degree of their neighbors, they are still locally dominated in a median-based sense. The juxtaposition of panels (a)–(c) therefore illustrates how majority-type median-based inequalities can be different from mean-based comparisons, even when both are derived from the same neighborhood degree data.
Examining the distribution of points across the four quadrants of panel (c) in Figs. 2 and 3 offers an immediate visual explanation for the observed values of $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$. In the ZKC network, most points fall in the first and third quadrants, where $k_{i}-k_{nn}(i)$ and $h_{i}-1/2$ share the same sign. Nodes that are disadvantaged in the mean-based sense tend also to be disadvantaged in the median-based sense, which explains why $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ take similarly large values. The AFB network presents a contrasting pattern: a substantial mass in the second quadrant corresponds to nodes with $k_{i}>k_{nn}(i)$ but $h_{i}<1/2$. These nodes have neighbors $\mathcal{N}_{i}$ whose distribution of degrees in $\mathcal{K}_{i}$ is sufficiently left-skewed that the mean is pulled below the median. This creates a mismatch between mean-based and median-based comparisons and results in $\phi_{\mathrm{global}}<1/2$ but $\phi_{\mathrm{local}}\approx 0.84$. In the hypothetical contrasting case of a right-skewed distribution where the nodes in the fourth quadrant of panel (c) is abundant, which results in many nodes with $k_{i}<k_{nn}(i)$ but $h_{i}>1/2$, it would be possible that $\phi_{\mathrm{global}}>1/2$ and $\phi_{\mathrm{local}}<1/2$.
Taken together, these examples show that the two classical FP inequalities and the two majority-type inequalities capture qualitatively different structural features. The FP remains a statement about network-level averages, whereas $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ describe majority-type neighborhood relations. Their systematic divergence across networks underscores the need to treat these quantities as distinct diagnostics rather than interchangeable formulations of “your friends have more friends than you.”
V Discussion and Outlook
The analysis in this work shows that a full account of how a node compares with its neighbors requires distinguishing four logically independent inequalities: the classical alter-based FP, expressed by $\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}\geq\langle k\rangle_{\mathrm{n}}$; the classical ego-based FP, expressed by $\langle k_{nn}\rangle_{\mathrm{n}}\geq\langle k\rangle_{\mathrm{n}}$; the global majority-type condition based on mean neighbor degree, $\phi_{\mathrm{global}}$ can be either $\geq 1/2$ or <1/2<1/2; and the local majority-type condition based on median neighbor degree, $\phi_{\mathrm{local}}$ can also be either $\geq 1/2$ or <1/2<1/2. The first two inequalities follow directly from the definitions of edge-based and node-based averages and hold for any network. The latter two, by contrast, are majority statements whose truth values are unconstrained by the classical FP. They capture structurally distinct neighborhood relations and should therefore be treated as different diagnostics rather than as reformulations of the FP itself.
The explicit construction in Fig. 1 demonstrates that both $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ can fall below one-half. The toy example highlights the fact that mean-based and median-based comparisons capture different local phenomena: a node may be dominated in the mean sense but not in the median sense, or vice versa. The empirical results in Table 1 reinforce this point. In the ZKC network, both $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ are high, indicating a broad consistency between mean-based and median-based local disadvantage. In contrast, the AFB network exhibits $\phi_{\mathrm{global}}<1/2$ yet $\phi_{\mathrm{local}}\approx 0.84$, showing that most teams are dominated by their neighbors in a median sense even though a majority have degree exceeding the mean of their neighbors’ degrees. The density plots in Figs. 2 and 3 illustrate these divergences visually: the signed contrast $k_{i}-k_{nn}(i)$ and the hub-centrality-based quantity $h_{i}$ need not be concomitant.
Overall, the results make clear that the classical FP concerns population-level averages, whereas the quantities $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ capture local majority relations. The quadrant structure in panel (c) of Figs. 2 and 3 shows that these majority-type inequalities depend sensitively on the joint distribution of $\left(h_{i},k_{nn}(i)\right)$ and on the shape of each node’s neighbor-degree distribution. In particular, left- or right-skewed degree distributions of neighboring nodes—when the mean is significantly different from the median—produce systematic mismatches between mean-based and median-based comparisons. These effects cannot be inferred from the classical FP alone and underscore that the four inequalities considered here diagnose structurally distinct forms of local advantage or disadvantage.
Several directions for future research follow naturally from this perspective. One avenue is the analytical characterization of $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$ in canonical random graph ensembles, including the configuration model and scale-free networks, with particular attention to the full distribution of $k_{nn}(i)$ rather than only its mean. A second direction is to investigate how assortativity [9], community structures [10], and core-periphery [11] or rich-club organization [12] shape both the four inequalities and the quadrant patterns observed in panel (c) of Figs. 2 and 3. Extending the analysis to temporal or adaptive networks [13] is equally promising, since growth, rewiring, or behavioral feedback can shift local degree contrasts in ways that classical FP averages cannot capture. Finally, applications in systems where local perception asymmetry has behavioral or functional consequences—such as social influence, information diffusion, infrastructure robustness, or neuronal circuits—may benefit from explicitly distinguishing mean-based and median-based notions of local disadvantage.
The hope is that viewing the FP alongside its mean-based and median-based local relatives will provide a flexible and interpretable language for describing how nodes perceive their structural position in complex networks. These quantities open the door to a more comprehensive theory of local comparison, extending the conceptual scope of the friendship paradox beyond its traditional mean-based foundations. In any case, the seemingly “paradoxical” situations in which a majority of nodes appear to “violate” the friendship paradox—what one might playfully call the friendship-paradox paradox (FPP)—are resolved within the framework developed here. The distinction between mean-based FP statements and majority-type, median-based inequalities clarifies why such phenomena arise and ensures that no genuine contradiction exists.
Epilogue: Learning from misunderstanding—classroom anecdote
The author first encountered the distinction between mean-based and majority-type comparisons in a particularly memorable moment while teaching a network science course. As part of a homework assignment, students were asked to examine the FP for the AFB network data [7, 8]. One student proudly reported that the FP does not hold in this network, pointing to the fact that $\phi_{\mathrm{global}}<1/2$: for a majority of teams, the mean degree of their neighbors is actually lower than their own degree. The conclusion might sound reasonable, but incorrect, as any reader who is reading this paper up to this point can easily tell.
What had happened was a misconception that the author suspects many readers may share initially: the student had implicitly replaced the FP’s comparison of averages with a comparison of majorities. The classical ego-based FP concerns Eq. (6), which compares network-wide means and indeed holds for the football network. The quantity $\phi_{\mathrm{global}}$, on the other hand, asks a completely different question: whether the majority of nodes individually satisfy $k_{i}<k_{nn}(i)$. There is no mathematical reason for this fraction to exceed one-half, and in the AFB network it clearly does not. This classroom episode served as a useful reminder that the colloquial phrasing of the FP can tempt readers into thinking about “most” instead of “average.” The mismatch revealed by the AFB network provides a concrete motivation for carefully separating mean-based FP statements from majority-type, median-oriented interpretations and for studying the latter in their own right.
Acknowledgements.
This research was supported by the National Research Foundation of Korea (NRF) under the grant RS-2021-NR061247. This research was supported by the National Research Foundation of Korea (NRF) under the grant RS-2021-NR061247. The author appreciates Krzysztof Sienicki for pointing out a subtle but important technical issue regarding the relationship between the condition $h_{i}<1/2$ and the neighborhood median, which helped improve the clarity and accuracy of the presentation.